# Online POMDP Planning with Anytime Deterministic Guarantees

### v1

@@ 3.1 Simplified observation space
Corollary 1.1
Notably, UDB does not require a complete recovery of the posterior branches to yield an optimal policy. Each action-value is bounded by a specific lower and upper bound, which can be represented as an interval enclosing the theoretical value. When the bound intervals of two candidate actions do not overlap, one can clearly discern which action is suboptimal, rendering its subtree redundant for further exploration. This distinction sets UDB apart from current state-of-the-art online POMDP algorithms. In those methods, any finite-time stopping condition fails to ensure optimality since the bounds used are either heuristic or probabilistic in nature.
3.2 Simplified State and Observation Spaces
In certain scenarios, the complete evaluation of posterior beliefs during the planning stage may pose significant computational challenges. To tackle this issue, we propose the use of a simplified state space in addition to the simplified observation space considered thus far. Specifically, we derive deterministic guarantees of the value function that allow for the selection of a subset from both the states and observations.
Theorem 2
Let $b_{0}$ and $\bar{b}_{0}$ be the theoretical and simplified belief states, respectively, at time $t=0$, and $T$ be the last time step of the POMDP. Let $V^{\pi}(b_{0})$ be the theoretical value function by following a policy $\pi$, and let $\bar{V}^{\pi}(\bar{b}_{0})$ be the simplified value function by following the same policy, as defined in (7). Then, for any policy $\pi$, the difference between the theoretical and simplified value functions is bounded as follows,
$\left|V^{\pi}(b_{0})-\bar{V}^{\pi}(\bar{b}_{0})\right|\leq\mathcal{R}_{\max}\left[1-\sum_{x}\bar{b}_{0}(x)\right]+\mathcal{R}_{\max}\sum_{\tau=1}^{T}\left[1-\overline{\mathbb{E}}_{z_{1:\tau}}\sum_{x}\overline{b}_{\tau}(x)\right]\triangleq\epsilon_{x,z}^{\pi}(b_{0}),$ (16)
Similar to Theorem 1, $\epsilon_{x,z}^{\pi}(b_{0})$ only contains probability values of instances from the simplified sets. However, this bound accounts for both the states and observations that are within the simplified spaces, $x,z\in\bar{\mathcal{X}},\bar{\mathcal{Z}}$, which makes it a viable to compute at planning time. In contrast to Theorem 1, this bound can only be calculated at the root since it relies on the knowledge of the actual probability value of the prior, $b_{0}(x)$, for states $x\in\bar{\mathcal{X}}$, which are only available at the root. Furthermore, since now the belief density function at different belief nodes is not required, we demonstrate in the next section that a belief update can be avoided completely, which makes it suitable to attach guarantees to state-of-the-art algorithms.
The process of maintaining an upper bound for action-value function follows similarly to the one presented in the observation-only simplification subsection 3.1,
$\epsilon_{x,z}^{\pi}(b_{t},a_{t})\triangleq\mathcal{R}_{\max}\left[1-\sum_{x_{t:t+1}}b(x_{t})\overline{\mathbb{P}}(x_{t+1}\mid x_{t},a_{t})\right]+$ (17)
$\mathcal{R}_{\max}\!\!\!\sum_{\tau=t+1}^{T}\!\!\left[1\!-\!\!\sum_{z_{t+1:{\tau}}}\!\sum_{x_{t:{\tau}}}b(x_{t})\overline{\mathbb{P}}(z_{t+1}\mid x_{t+1})\overline{\mathbb{P}}(x_{t+1}\mid x_{t},a_{t})\!\!\!\prod_{k=t+2}^{\tau}\!\!\overline{\mathbb{P}}(z_{k}\mid x_{k})\overline{\mathbb{P}}(x_{k}\mid x_{k-1},\pi_{k-1})\right].$
In $\epsilon_{x,z}^{\pi}(b_{t},a_{t})$, the subscript of $(x,z)$ differentiates the bound of the state- and observation-simplification from the observation-only simplification bound.
Computing the bounds in practice is straightforward and can be done iteratively, which fits many online planning algorithms. As it requires the summation over probability values for every considered trajectory, ${x_{0},a_{0},x_{1},z_{1},...,a_{t-1},x_{t},z_{t}}$, as shown in equations (17) and (11). In the subsequent section, we provide further details and illustrate how the bounds derived from the state-observation simplification can be used to derive deterministic guarantees on policies obtained from existing state-of-the-art algorithms.
4 Methods
Algorithm 1 Algorithm-$\mathcal{A}$:
0:   function Search 1:  while time permits do 2:     Generate states $x$ from $b_{0}$. 3:     $\tau_{0}\xleftarrow{}x$ 4:     ¯$\bar{\mathbb{P}}_{0}\xleftarrow{}b(x_{0}\mid h_{0})$ 5:     if $\tau_{0}\notin\tau(h)$ then 6:        ¯$\bar{\mathbb{P}}(h)\xleftarrow{}\bar{\mathbb{P}}(h)+\bar{\mathbb{P}}_{0}$ 7:     end if 8:     $\textsc{Simulate}(h,D,\tau_{0},\bar{\mathbb{P}}_{0})$. 9:  end while 10:  return   0:   function fwdUpdate($ha,haz$, $\tau_{d}$, ¯$\bar{\mathbb{P}}_{\tau}$, $x^{\prime}$) 1:  if $\tau_{d}\notin\tau(ha)$ then 2:     $\tau(ha)\xleftarrow{}\tau(ha)\cup\{\tau_{d}\}$ 3:     $\bar{R}(ha)\xleftarrow{}\bar{R}(ha)+\bar{\mathbb{P}}_{\tau}\cdot r(x,a)$ 4:  end if 5:  $\tau_{d}\xleftarrow{}\tau_{d}\cup\{x^{\prime}\}$ 6:  ¯$\bar{\mathbb{P}}_{\tau}\xleftarrow{}\bar{\mathbb{P}}_{\tau}\cdot Z_{z\mid x^{\prime}}\cdot T_{x^{\prime}\mid x,a}$ 7:  if $\tau_{d}\notin\tau(haz)$ then 8:     ¯$\bar{\mathbb{P}}(haz)\xleftarrow{}\bar{\mathbb{P}}(haz)+\bar{\mathbb{P}}_{\tau}$ 9:     $\tau(haz)\xleftarrow{}\tau(haz)\cup\{\tau_{d}\}$ 10:  end if 11:  return
0:   function $\textsc{Simulate}(h,d,\tau_{d},\bar{\mathbb{P}}_{d})$ 1:  if $d=0$ then 2:     return  $\delta(h)\xleftarrow{}0$ 3:  end if 4:  Select action $a$. 5:  Generate next states and observations, $x^{\prime},z$. 6:  $\tau_{d},\ \bar{\mathbb{P}}_{\tau}\xleftarrow{}$fwdUpdate($ha,haz,\tau_{d},\bar{\mathbb{P}}_{\tau},x^{\prime}$) 7:  Select next observation $z$. 8:  Simulate$(haz,d-1,\tau_{d},\bar{\mathbb{P}}_{\tau})$ 9:  bwdUpdate($h,ha,d$) 10:  return   0:   function bwdUpdate($h,ha,d$) 1:  $\epsilon(ha)=\gamma^{D-d}V_{\max}(\bar{\mathbb{P}}(h)-\bar{\mathbb{P}}(ha))+\gamma^{D-d-1}\cdot V_{\max}(\bar{\mathbb{P}}(ha)-\sum\limits_{z\mid ha}\bar{\mathbb{P}}(haz))$ 2:  $U(ha)\!=\!\bar{R}(ha)+\gamma\!\sum_{z\mid ha}\!U(haz)+\epsilon(ha)$ 3:  $L(ha)\!=\!\bar{R}(ha)+\gamma\!\sum_{z\mid ha}\!L(haz)-\epsilon(ha)$ 4:  $U(h)\xleftarrow{}\max_{a^{\prime}}\{U(ha^{\prime})\}$ 5:  $L(h)\xleftarrow{}\max_{a^{\prime}}\{L(ha^{\prime})\}$ 6:  return
In this section we aim to describe how to fit our bounds to a general structure algorithm, named $\textsc{Algorithm}-\mathcal{A}$, which serves as an abstraction to many existing algorithms. To compute the deterministic bounds, we utilize Bellman’s update and optimality criteria. This approach naturally fits dynamic programming approaches such as DESPOT (Ye et al., 2017) and AdaOPS (Wu et al., 2021). However, it may also be attached with algorithms that rely on Monte-Carlo estimation, such as POMCP (Silver and Veness, 2010), by viewing the search tree as a policy tree.
[…]

### v4

@@ 9.2 Lemma 1
Proof.
In the following, we prove by induction that applying the Bellman optimality operator on upper bounds to the value function in finite-horizon POMDPs will result in an upper bound on the optimal value function. The notations are the same as the ones presented in the main body of the paper. We restate some of the definitions from the paper for convenience.
The policy $\pi_{t}(b_{t})$ determined by applying Bellman optimality at belief $b_{t}$, i.e.,
$\pi_{t}(b_{t})=\arg\max_{a_{t}\in\mathcal{A}}[\bar{Q}^{\pi}(b_{t},a_{t})+\epsilon_{z}^{\pi}(b_{t},a_{t})].$ (57)
As it will be needed in the following proof, we also define the value of a belief which includes in its history at least one observation out of the simplified set, e.g. $H_{t}=\{a_{0},z_{1},\dotsc,z_{k}\notin\bar{\mathcal{Z}},\dotsc,z_{t}\}$ as being equal to zero. Explicitly,
$\bar{V}_{t}^{\pi}(\mathbb{P}(x_{t}\mid a_{0},z_{1},\dotsc,z_{k}\notin\bar{\mathcal{Z}},\dotsc,z_{t}))\equiv 0\ \ \forall k\in[1,t].$ (58)
We also use the following simple bound,
$V_{t,\max}\triangleq\mathcal{R}_{\max}\cdot(T-t-1)$ (59)
Base case $(t=T)$ At the final time step $T$, for each belief we set the value function to be equal to the reward value at that belief state, $b_{T}$ and taking the action that maximizes the immediate reward,
$\textsc{UDB}^{\pi}(b_{T})=\max_{a_{T}}\{r(b_{T},a_{T})+\epsilon_{z}(b_{T},a_{T})\}\equiv\arg\max_{a_{T}}\{r(b_{T},a_{T})\}$ (60)
which provides an upper bound for the optimal value function for the final time step, $V_{T}^{\star}(b_{T})\leq\textsc{UDB}^{\pi}(b_{T})$.
Induction hypothesis - Assume that for a given time step, $t$, for all belief states the following holds,
$V_{t}^{\star}(b_{t})\leq\textsc{UDB}^{\pi}(b_{t}).$ (61)
Induction step - We will show that the hypothesis holds for time step $t-1$. By the induction hypothesis,
$V_{t}^{\star}(b_{t})\leq\textsc{UDB}^{\pi}(b_{t})\ \ \forall b_{t},$ (62)
thus,
$Q^{\star}(b_{t-1},a_{t-1})=r(b_{t-1},a_{t-1})+\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)V_{t}^{\star}(b(z_{t}))$ (63)
$\leq r(b_{t-1},a_{t-1})+\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\textsc{UDB}^{\pi}(b(z_{t}))$ (64)
$=r(b_{t-1},a_{t-1})+\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\left[\bar{V}_{t}^{\pi}(b_{t})+\epsilon_{z}^{\pi}(b_{t})\right].$ (65)
For the following transition, we make use of lemma 5,
$=r(b_{t-1},a_{t-1})+\bar{\mathbb{E}}_{z_{t}\mid b_{t-1},a_{t-1}}\left[\bar{V}_{t}^{\pi}(b_{t})\right]+\epsilon_{z}^{\pi}(b_{t-1},a_{t-1})$ (66)
$\equiv\textsc{UDB}^{\pi}(b_{t-1},a_{t-1}).$ (67)
Therefore, under the induction hypothesis, $Q_{t-1}^{\star}(b_{t-1},a_{t-1})\leq\textsc{UDB}^{\pi}(b_{t-1},a_{t-1})$. Taking the maximum over all actions $a_{t}$,
$\textsc{UDB}^{\pi}(b_{t-1})=\max_{a_{t-1}\in\mathcal{A}}\left\{\textsc{UDB}^{\pi}(b_{t-1},a_{t-1})\right\}$ (68)
$\geq\max_{a_{t-1}\in\mathcal{A}}\left\{Q_{t-1}^{\star}(b_{t-1},a_{t-1})\right\}=V_{t-1}^{\star}(b_{t-1}),$
which completes the induction step and the required proof. ∎
Lemma 5.
Let $b_{t}$ denote a belief state and $\pi_{t}$ a policy at time $t$. Let ¯$\bar{\mathbb{P}}(z_{t}\mid x_{t})$ be the simplified observation model which represents the likelihood of observing $z_{t}$ given $x_{t}$. Then, the following terms are equivalent,
$\mathbb{E}_{z_{t}}\left[\bar{V}_{t}^{\pi}(b_{t})+\epsilon_{z}^{\pi}(b_{t})\right]=\bar{\mathbb{E}}_{z_{t}}\left[\bar{V}_{t}^{\pi}(b_{t})\right]+\epsilon_{z}^{\pi}(b_{t-1},a_{t-1})$ (69)
Proof.
$\mathbb{E}_{z_{t}}\left[\bar{V}_{t}^{\pi}(b_{t})+\epsilon_{z}^{\pi}(b_{t})\right]=$ (70)
$\mathbb{E}_{z_{t}}\left[\bar{V}_{t}^{\pi}(b_{t})\right]+\mathbb{E}_{z_{t}}\left[\mathcal{R}_{\max}\sum_{\tau=t+1}^{T}\left[1-\sum_{z_{t+1:\tau}}\sum_{x_{t:\tau}}b_{t}\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]\right]$ (71)
focusing on the second summand,
$\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\mathcal{R}_{\max}\sum_{\tau=t+1}^{T}\left[1-\sum_{z_{t+1:\tau}}\sum_{x_{t:\tau}}b_{t}\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]$ (72)
$=\mathcal{R}_{\max}\sum_{\tau=t+1}^{T}\left[1-\sum_{z_{t}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\sum_{z_{t+1:\tau}}\sum_{x_{t:\tau}}b(x_{t})\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]$ (73)
by marginalizing over $x_{t-1}$,
$=\mathcal{R}_{\max}\sum_{\tau=t+1}^{T}[1-\sum_{z_{t}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\sum_{z_{t+1:\tau}}\sum_{x_{t-1:\tau}}\frac{\bar{\mathbb{P}}(z_{t}\mid x_{t})\mathbb{P}(x_{t}\mid x_{t-1},\pi_{t-1})b(x_{t-1})}{\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)}\cdot$ (74)
$\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})]$
canceling out the denominator,
$=\mathcal{R}_{\max}\sum_{\tau=t+1}^{T}[1-\sum_{z_{t:\tau}}\sum_{x_{t-1:\tau}}\bar{\mathbb{P}}(z_{t}\mid x_{t})\mathbb{P}(x_{t}\mid x_{t-1},a_{t-1})b(x_{t-1})\cdot$ (75)
$\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})]\equiv\epsilon_{z}^{\pi}(b_{t-1},a_{t-1})$
it is left to show that $\mathbb{E}_{z_{t}\mid b_{t-1},a_{t-1}}\left[\bar{V}_{t}^{\pi}(b_{t})\right]=\bar{\mathbb{E}}_{z_{t}\mid b_{t-1},a_{t-1}}\left[\bar{V}_{t}^{\pi}(b_{t})\right]$. By the definition of a value function of a belief not included in the simplified set, we have that,
$\mathbb{E}_{z_{t}\mid b_{t-1},a_{t-1}}\left[\bar{V}_{t}^{\pi}(b_{t})\right]$ $=\sum_{z_{t}\in\mathcal{Z}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\bar{V}_{t}^{\pi}(b_{t})$ (76)
$=\sum_{z_{t}\in\bar{\mathcal{Z}}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\bar{V}_{t}^{\pi}(b_{t})+\sum_{z_{t}\in\mathcal{Z}\backslash\bar{\mathcal{Z}}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\bar{V}_{t}^{\pi}(b_{t})$ (77)
$=\sum_{z_{t}\in\bar{\mathcal{Z}}}\bar{\mathbb{P}}\left(z_{t}\mid H_{t}^{-}\right)\cdot\bar{V}_{t}^{\pi}(b_{t})+\sum_{z_{t}\in\mathcal{Z}\backslash\bar{\mathcal{Z}}}\mathbb{P}\left(z_{t}\mid H_{t}^{-}\right)\cdot 0$ (78)
$=\bar{\mathbb{E}}_{z_{t}\mid b_{t-1},a_{t-1}}\left[\bar{V}_{t}^{\pi}(b_{t})\right],$ (79)
which concludes the derivation. ∎
9.3 Corollary 1.1
We restate the definition of UDB exploration criteria,
Corollary 3.1.
Using Lemma 1 and the exploration criteria defined in (17) guarantees convergence to the optimal value function.
Proof.
Let us define a sequence of bounds, $\textsc{UDB}_{n}^{\pi}(b_{t})$ and a corresponding difference value between $\textsc{UDB}_{n}$ and the simplified value function,
$\textsc{UDB}_{n}^{\pi}(b_{t})-\bar{V}_{n}^{\pi}(b_{t})=\epsilon_{n,z}^{\pi}(b_{t}),$ (81)
where $n\in[0,\left|\mathcal{Z}\right|]$ corresponds to the number of unique observation instances within the simplified observation set, $\bar{\mathcal{Z}}_{n}$, and $\left|\mathcal{Z}\right|$ denotes the cardinality of the complete observation space. Additionally, for the clarity of the proof and notations, assume that by construction the simplified set is chosen such that $\bar{\mathcal{Z}}_{n}(H_{t})\equiv\bar{\mathcal{Z}}_{n}$ remains identical for all time steps $t$ and history sequences, $H_{t}$ given $n$. By the definition of $\epsilon_{n,z}^{\pi}(b_{t})$,
$\epsilon_{n,z}^{\pi}(b_{t})=\mathcal{R}_{\max}\!\!\!\sum_{\tau=t+1}^{T}\!\!\left[1\!-\!\!\!\sum_{z_{t+1:{\tau}}\in\bar{\mathcal{Z}}_{n}}\sum_{x_{t:{\tau}}}b(x_{t})\!\!\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right],$ (82)
we have that $\epsilon_{n,z}^{\pi}(b_{t})\rightarrow 0$ as $n\rightarrow\left|\mathcal{Z}\right|$, since
$\sum_{z_{t+1:{\tau}}\in\bar{\mathcal{Z}}_{n}}\sum_{x_{t:{\tau}}}b(x_{t})\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\rightarrow 1$ (83)
as more unique observation elements are added to the simplified observation space, $\bar{\mathcal{Z}}_{n}$, eventually recovering the entire support of the discrete observation distribution.
From lemma 1 we have that, for all $n\in[0,\left|\mathcal{Z}\right|]$ the following holds,
$V^{\pi*}(b_{t})\leq\textsc{UDB}_{n}^{\pi}(b_{t})=\bar{V}_{n}^{\pi}(b_{t})+\epsilon_{n,z}^{\pi}(b_{t}).$ (84)
Additionally, from theorem 1 we have that,
$\left|V^{\pi}(b_{t})-\bar{V}_{n}^{\pi}(b_{t})\right|\leq\epsilon_{n,z}^{\pi}(b_{t}),$ (85)
for any policy $\pi$ and subset $\bar{\mathcal{Z}}_{n}\subseteq\mathcal{Z}$, thus,
$\bar{V}_{n}^{\pi}(b_{t})-\epsilon_{n,z}^{\pi}(b_{t})\leq V^{\pi}(b_{t})\leq V^{\pi*}(b_{t})\leq\bar{V}_{n}^{\pi}(b_{t})+\epsilon_{n,z}^{\pi}(b_{t}).$ (86)
Since $\epsilon_{n,z}^{\pi}(b_{t})\rightarrow 0$ as $n\rightarrow\left|\mathcal{Z}\right|$, and $\left|\mathcal{Z}\right|$ is finite, it is guaranteed that $\textsc{UDB}_{n}^{\pi}(b_{t})\xrightarrow{n\rightarrow\left|\mathcal{Z}\right|}V^{\pi*}(b_{t})$ which completes our proof. ∎
Moreover, depending on the algorithm implementation, the number of iterations can be finite (e.g. by directly choosing actions and observations to minimize the bound). A stopping criteria can also be verified by calculating the difference between the upper and lower bounds. The optimal solution is obtained once the upper bound equals the lower bound.
9.4 Theorem 2
Theorem 4.
$\left|V^{\pi}(b_{t})\!-\!\bar{V}^{\pi}(b_{t})\right|\leq\!\mathcal{R}_{\max}\!\!\!\sum_{\tau=t+1}^{T}\!\!\left[1\!-\!\!\!\sum_{z_{t+1:{\tau}}}\sum_{x_{t:{\tau}}}b(x_{t})\!\!\prod_{k=t+1}^{\tau}\bar{\mathbb{P}}(z_{k}\mid x_{k})\mathbb{P}(x_{k}\mid x_{k-1},\pi_{k-1})\right]\triangleq\epsilon^{\pi}(b_{t}).$ (87)
Recall that we define $\tau_{t}=\{x_{0},a_{0},z_{1},x_{1},a_{1},\dotsc,a_{T-1},x_{t},z_{t}\}$. Then the value function is defined as,
$V^{\pi}(b_{0})=\sum_{\tau_{T}}\mathbb{P}^{\pi}(\tau_{T})\left[\sum_{t=0}^{T}r(x_{t},a_{t})\right]$ (88)
applying chain rule and rearranging terms,
$=\sum_{\tau_{T}}\mathbb{P}^{\pi}(x_{1:T},z_{1:T},a_{1:T}\mid\tau_{0})\mathbb{P}^{\pi}(\tau_{0})\left[\sum_{t=0}^{T}r(x_{t},a_{t})\right]$ (89)
$=\sum_{\tau_{0}}\mathbb{P}^{\pi}(\tau_{0})\sum_{x_{1:T},z_{1:T},a_{1:T}}\mathbb{P}^{\pi}(x_{1:T},z_{1:T},a_{1:T}\mid\tau_{0})\left[\sum_{t=0}^{T}r(x_{t},a_{t})\right]$ (90)
$=\sum_{\tau_{0}}\mathbb{P}^{\pi}(\tau_{0})\left[r(x_{0},a_{0})+\sum_{x_{1:T},z_{1:T},a_{1:T}}\mathbb{P}^{\pi}(x_{1:T},z_{1:T},a_{1:T}\mid\tau_{0})\left[\sum_{t=1}^{T}r(x_{t},a_{t})\right]\right]$ (91)
nullifying instances of the complete probability distribution, $\mathbb{P}^{\pi}(\cdot)$, is denoted as a simplified distribution, ¯$\bar{\mathbb{P}}^{\pi}(\cdot)$. We can then split and bound from above the value function, such that the simplified value function consideres only a subset of the trajectories at time $t=0$,
$\leq\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\left[r(x_{0},a_{0})+\sum_{x_{1:T},z_{1:T},a_{1:T}}\mathbb{P}^{\pi}(x_{1:T},z_{1:T},a_{1:T}\mid\tau_{0})\left[\sum_{t=1}^{T}r(x_{t},a_{t})\right]\right]$ (92)
$+\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]\mathcal{V}_{max,0}$ (93)
We then apply similar steps on the next time step, $t=1$,
$=$ $\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\left[r(x_{0},a_{0})+\sum_{x_{1:T},z_{1:T},a_{1:T}}\mathbb{P}^{\pi}(x_{2:T},z_{2:T},a_{2:T}\mid\tau_{1})\mathbb{P}^{\pi}(x_{1},z_{1},a_{1}\mid\tau_{0})\left[\sum_{t=1}^{T}r(x_{t},a_{t})\right]\right]$ (94)
$+\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]\mathcal{V}_{max,0}$ (95)
$=$ $\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\Biggl{[}r(x_{0},a_{0})$ (96)
$+\sum_{x_{1},z_{1},a_{1}}\mathbb{P}^{\pi}(x_{1},z_{1},a_{1}\mid\tau_{0})\sum_{x_{2:T},z_{2:T},a_{2:T}}\mathbb{P}^{\pi}(x_{2:T},z_{2:T},a_{2:T}\mid\tau_{1})\left[\sum_{t=1}^{T}r(x_{t},a_{t})\right]\Biggr{]}$
$+\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]\mathcal{V}_{max,0}$ (97)
$=$ $\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\Biggl{[}r(x_{0},a_{0})$ (98)
$+\sum_{x_{1},z_{1},a_{1}}\mathbb{P}^{\pi}(x_{1},z_{1},a_{1}\mid\tau_{0})\Biggl{[}r(x_{1},a_{1})$
$+\sum_{x_{2:T},z_{2:T},a_{2:T}}\mathbb{P}^{\pi}(x_{2:T},z_{2:T},a_{2:T}\mid\tau_{1})\left[\sum_{t=2}^{T}r(x_{t},a_{t})\right]\Biggr{]}\Biggr{]}$
$+\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]\mathcal{V}_{max,0}$
$\leq$ $\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\Biggl{[}r(x_{0},a_{0})$ (99)
$+\sum_{x_{1},z_{1},a_{1}}\bar{\mathbb{P}}^{\pi}(x_{1},z_{1},a_{1}\mid\tau_{0})\Biggl{[}r(x_{1},a_{1})$
$+\sum_{x_{2:T},z_{2:T},a_{2:T}}\mathbb{P}^{\pi}(x_{2:T},z_{2:T},a_{2:T}\mid\tau_{1})\left[\sum_{t=2}^{T}r(x_{t},a_{t})\right]\Biggr{]}\Biggr{]}$
$+\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\left[1-\sum_{x_{1},z_{1},a_{1}}\bar{\mathbb{P}}^{\pi}(x_{1},z_{1},a_{1}\mid\tau_{0})\right]\mathcal{V}_{max,1}+\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]\mathcal{V}_{max,0}$ (100)
which results in,
$=\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\Biggl{[}r(x_{0},a_{0})$ (101)
$+\sum_{x_{1},z_{1},a_{1}}\bar{\mathbb{P}}^{\pi}(x_{1},z_{1},a_{1}\mid x_{0},a_{0})\left[r(x_{1},a_{1})+\sum_{x_{2:T},z_{2:T},a_{2:T}}\mathbb{P}^{\pi}(x_{2:T},z_{2:T},a_{2:T}\mid\tau_{1})\left[\sum_{t=2}^{T}r(x_{t},a_{t})\right]\right]\Biggr{]}$
$+\left[\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})-\sum_{\tau_{1}}\bar{\mathbb{P}}^{\pi}(\tau_{1})\right]\mathcal{V}_{max,1}+\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]\mathcal{V}_{max,0}$ (102)
Performing the same steps iteratively up to time $t=T$, yields the desired outcome,
$V^{\pi}(b_{0})\leq\sum_{t=0}^{T}\sum_{\tau_{t}}\bar{\mathbb{P}}^{\pi}(\tau_{t})r(x_{t},a_{t})+\mathcal{V}_{max,0}\left[1-\sum_{\tau_{0}}\bar{\mathbb{P}}^{\pi}(\tau_{0})\right]+\sum_{t=0}^{T-1}\mathcal{V}_{max,t+1}\left[\sum_{\tau_{t}}\bar{\mathbb{P}}^{\pi}(\tau_{t})-\sum_{\tau_{t+1}}\bar{\mathbb{P}}^{\pi}(\tau_{t+1})\right]$ (103)
9.5 Optimality Guarantees
Lemma 6.
Let $\mathcal{A}$ be the set of actions and $\mathcal{U}_{0}^{\star}(H_{t})$, $\mathcal{L}_{0}^{\star}(H_{t})$ be the upper and lower bounds of node $H_{t}$ chosen according to,
$\mathcal{U}^{\star}_{0}(H_{t})\triangleq\sum_{\tau_{t}\in\mathcal{T}(H_{t})}\bar{\mathbb{P}}(\tau_{t})\left[r(x_{t},a_{t})+\mathcal{V}_{\max,t}\right]+\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},a_{t})}\left[\mathcal{U}^{\star}_{0}(H_{t+1})-\sum_{\tau_{t+1}\in\mathcal{T}(H_{t+1})}\!\!\!\!\!\!\!\!\!\bar{\mathbb{P}}(\tau_{t+1})\mathcal{V}_{\max,t}\right]$ (104)
$\mathcal{L}^{\star}_{0}(H_{t})\triangleq\sum_{\tau_{t}\in\mathcal{T}(H_{t})}\bar{\mathbb{P}}(\tau_{t})\left[r(x_{t},a_{t})+\mathcal{V}_{\min,t}\right]+\sum_{z_{t+1}\in\bar{\mathcal{Z}}(H_{t},a_{t})}\left[\mathcal{L}^{\star}_{0}(H_{t+1})-\sum_{\tau_{t+1}\in\mathcal{T}(H_{t+1})}\!\!\!\!\!\!\!\!\!\bar{\mathbb{P}}(\tau_{t+1})\mathcal{V}_{\min,t}\right]$ (105)
and,
$\mathcal{U}_{0}^{\star}(H_{T})\triangleq\sum_{\tau_{T}\in\mathcal{T}(H_{T})}\bar{\mathbb{P}}(\tau_{T})r(x_{T}),$ $\mathcal{L}_{0}^{\star}(H_{T})\triangleq\sum_{\tau_{T}\in\mathcal{T}(H_{T})}\bar{\mathbb{P}}(\tau_{T})r(x_{T}).$ (106)
Then, the optimal root-value is bounded by,
$\mathcal{L}_{0}^{\star}(H_{0})\leq V^{\pi^{*}}(H_{0})\leq\mathcal{U}_{0}^{\star}(H_{0}).$ (107)
Proof.
We wish to show that $\mathcal{L}_{0}^{\star}(H_{0})\leq V^{\pi^{*}}(b_{0})\leq\mathcal{U}_{0}^{\star}(H_{0})$. We derive a proof for one side of the inequality, while the other follows similarly. First note that,
$V^{\pi^{*}}(b_{0})\leq\mathcal{U}_{0}^{\pi^{*}}(H_{0})\leq\underset{\pi\in\Pi}{\max}\mathcal{U}_{0}^{\pi}(H_{0})$ (108)
where the first inequality is due to Theorem 2, and the second inequality is true by definition. However, the claim in Lemma 2 is a recursive claim, while the bound provided in Theorem 2 only holds with respect to the root. Thus, for completeness, we also need to show that the best action can be chosen recursively, even though the bound is ‘partial‘ in different parts of the tree.
$\underset{\pi_{0:T}\in\Pi}{\max}\mathcal{U}_{0}^{\pi}(H_{0})$
$=\underset{\pi_{0:T}\in\Pi}{\max}\sum_{\tau_{0}\in\mathcal{T}(H_{0})}\overline{\mathbb{P}}(\tau_{0})[r(x_{0},\pi_{0})+\mathcal{V}_{max,0}]+\sum_{z_{1}\in\overline{\mathcal{Z}}(H_{0},\pi_{0})}\left[\mathcal{U}_{0}^{\pi}(H_{1})-\sum_{\tau_{1}\in\mathcal{T}(H_{1})}\overline{\mathbb{P}}(\tau_{1})\mathcal{V}_{max,0}\right]$
[…]